OpenStudy (mylahb):

Ron is selling items at a craft show. The first hour his sales are $20. His sales grow by 10% each hour. What are Ron's total sales for the first 5 hour? A. $101.00 B. $102.02 C. $122.10

OpenStudy (anonymous):

The answer is C. 1h: 20 (10%= 2) 2h: 20 + 2 = 22 (10%= 2.2) 3h: 22 + 2.2 = 24.2 (10%= 2.42) 4h: 24.2 + 2.42 = 26.62 (10% = 2.66) 5h: 26.62 + 2.66 = 29. 28. 1h + 2h+ 3h+ 4h+ 5h= $122.10. I hope I could help, let me know if I was right, please. :)